# Computable Følner sequences of amenable groups

### v1

@@ 5 Metric groups and amenability
Locally compact groups form the basic area of classical amenability theory. Thus analysis of computability aspects of amenability in the case of metric locally compact groups is a natural challange. On the other hand motivated by strong progress made in the last two decates in the general case of amenable topological groups (see [15] [21], [30], [31], [32]), we study the subject without the restriction of local compactness.
@@ 5.1 Computable presentations of metric groups
We usually assume that a metric group is considered with a right-invariant metric $\leq 1$. This is a more general case compared to papers mentioned above, where bi-invariantness is usually assumed.
@@
  • If a homomorphism $\nu:(\mathbb{N},\star,^{-1},1)\to G$ is a computably enumerable presentation of $G$ and the sets $\mathsf{Wrd}^{=}$ and $\mathsf{Wrd}^{<}$ are computable, then we say that $\nu$ is a computable presentation and the group $(G,d,\nu)$ is computable.
Note that if $(G,d,\nu)$ is a numbered metric group such that the sets $\mathsf{Wrd}^{=}_{\nu}$ and $\mathsf{Wrd}^{<}_{\nu}$ are computably enumerable, then the group $(G,d,\nu_{1})$ built after Definition 5.1 is computably enumerable too. Indeed, since any word $u(\nu_{1}(m_{1}),\ldots,\nu_{1}(m_{s}))$ coincides with some $w(\nu(n_{1}),\ldots,\nu(n_{s}))$ and the latter one can be found in a computable way, any enumeration of the sets $\mathsf{Wrd}^{=}$ and $\mathsf{Wrd}^{<}$ with respect to $\nu$ effectively determines an enumeration of the corresponding set with respect to $\nu_{1}$.
@@
From now on we will consider computably enumerable groups such that $\nu$ is a homomorphism from the computable enumerated group $(\mathbb{N},*,^{-1},1)$ and the image of the numbering is a dense subgroup.
@@
  • If $(G,\nu)$ is computable in the sense of Section 2.1, then $(G,d,\nu)$ is a computable presentation in the sense of Section 2.4.
@@ 5.2 Amenability and effective amenability
The following notion will be convenient too.
Definition 5.7.
Given $\varepsilon$ and $g\in G$ we call an injective map $\phi:F\to gF$ an $\varepsilon$-matching for $gF$ if for every $g^{\prime}\in F$, $d(g^{\prime},\phi(g^{\prime}))<\varepsilon$ and the size of $\mathsf{dom}(\phi)$ is maximal for this condition.
@@
It is worth noting here that when an open neighbourhood $V$ contains $U$ the number $\mu(F,gF,U)$ does not exceed $\mu(F,gF,V)$. In particular in the formulation above we may consider neighbourhoods $U$ from a fixed base of identity neighbourhoods. For example in the case of a right-invariant metric group $(G,d)$ we may take all $U$ in the form of metric balls $B_{<q}=\{x:d(1,x)<q\}$, $q\in\mathbb{Q}\cap(0,1)$. It is also clear that we can restrict all $\theta$ by rational ones. From now on we work in this case.
@@ 5.3 Effective amenability of computable metric groups
(1) The assumptions of the first sentence of this proposition imply that the set $\mathsf{Wrd}^{=}$ is computable. To see this use the assumption that the multiplication * on $\mathsf{N}$ is a computable function.
@@
Let us fix an enumeration of all quadrangles of the form $(m,n,D,F)$ where $m,n\in\mathbb{N}$ and $D,F\subset_{fin}\mathbb{N}$. The following procedure, denoted below by $\hat{\vartheta}(m,n,D,F)$, determines quadrangles satisfying the condition of the proposition.
For an input $(m,n,D,F)$ let $F_{0}\subseteq F\cup D*F$ be a set representing the $\mathsf{T}_{=}$-classes in $F\cup E*F$. Let us fix an enumeration of the set
@@
If this is not the case we make the next step. Note that the number $\mu(\nu(F_{0}\cap F),\nu(g)\nu(F_{0}\cap F),B_{<\frac{1}{m}}(1))$ does nor decrease. Indeed, since the distances are becoming smaller, new matchings can occur, but the matchings which were already found, can only increase to larger ones. We stop when the above inequality holds. Note that if in the numbered metric group $(G,d,\nu)$ the condition
@@
Fix $n_{1},n_{2}$. Let $D$ be the set $\{n_{1},n_{2}\}$. We apply the algorithm for (ii) to some $(\ell,m,n,D)$ where $\frac{2}{\ell}\leq\varepsilon$ , $m>\frac{4}{\varepsilon}$ and $\frac{2}{3}\leq\frac{n-1}{n}$. Let $F\subset\mathbb{N}$ be a set which is the output of the algorithm and let $\Sigma^{\ell}_{F}=\{(f,f^{\prime},q)\,|\,f,f^{\prime}\in F,\,q\in\mathbb{Q}^{+},d(\nu(f),\nu(f^{\prime}))\in[q,q+\frac{1}{\ell})\}$.
@@
As in Section 4 an effective Følner sequence of the group $(G,d,\nu)$ is an effective sequence $(n_{j})_{j\in\mathbb{N}}$ where each $n_{j}$ is a Gödel number of some $F_{j}$, such that $(F_{j})_{j\in\mathbb{N}}$ is a Følner sequence.
@@
$(\phi(m,y)$ $\text{ is a total function})\wedge(\forall g\in\nu(\mathbb{N}))(\forall n)(\exists l)(\forall k)\Big{(}k>l\wedge(\phi(m,k)=f)$

### v2

@@ 5 Metric groups and amenability
Locally compact groups form the basic area of classical amenability theory. Thus analysis of computability aspects of amenability in the case of metric locally compact groups is a natural challenge. On the other hand motivated by strong progress made in the last two decates in the general case of amenable topological groups (see [14] [20], [29], [30], [31]), we study the subject without the restriction of local compactness.
@@ 5.1 Computable presentations of metric groups
We usually assume that a metric group is taken with a right-invariant metric $\leq 1$. This is a more general case compared to papers mentioned above, where bi-invariantness of the metric is assumed.
@@
  • If a homomorphism $\nu:(\mathbb{N},\star,^{-1},1)\to G$ is a computably enumerable presentation of $G$ and the sets $\mathsf{Wrd}^{=}_{\nu}$ and $\mathsf{Wrd}^{<}_{\nu}$ are computable, then we say that $\nu$ is a computable presentation and the group $(G,d,\nu)$ is computable.
Note that if $(G,d,\nu)$ is a numbered metric group such that the sets $\mathsf{Wrd}^{=}_{\nu}$ and $\mathsf{Wrd}^{<}_{\nu}$ are computably enumerable, then the group $(G,d,\nu_{1})$ built after Definition 5.1 is computably enumerable too. Indeed, since any word $u(\nu_{1}(m_{1}),\ldots,\nu_{1}(m_{s}))$ coincides with some $w(\nu(n_{1}),\ldots,\nu(n_{s}))$ and the latter one can be found in a computable way, any enumeration of the sets $\mathsf{Wrd}^{=}_{\nu}$ and $\mathsf{Wrd}^{<}_{\nu}$ with respect to $\nu$ effectively determines an enumeration of the corresponding set with respect to $\nu_{1}$.
@@
$\bullet$ From now on we will consider computably enumerable groups under such presentations that $\nu$ is a homomorphism from an enumerated group $(\mathbb{N},\star,^{-1},1)$ with computable operations and the image of the numbering is a dense subgroup.
@@
  • If $(G,\nu)$ is computable in the sense of Section 2.1, then $(G,d,\nu)$ is a computable presentation in the sense of Definition 5.2.
@@ 5.2 Amenability and effective amenability
It is worth noting here that when an open neighbourhood $V$ contains $U$ the number $\mu(F,gF,U)$ does not exceed $\mu(F,gF,V)$. In particular, in the formulation above we may consider neighbourhoods $U$ from a fixed base of identity neighbourhoods. For example in the case of a right-invariant metric group $(G,d)$ we may take all $U$ in the form of metric balls $B_{<q}=\{x:d(1,x)<q\}$, $q\in\mathbb{Q}\cap(0,1)$. It is also clear that we can restrict all $\theta$ by rational ones. From now on we work in this case.
@@
From the point of view of the terminology of Sections 2 and 3, it is more natural to include into this family sets $\nu(F)$ instead of $F$. However this would be slightly incovenient below.
@@ 5.3 Effective amenability of computable metric groups
(1) The assumptions of the first sentence of this proposition imply that the set $\mathsf{Wrd}^{=}_{\nu}$ (Definition 5.2) is computable. To see this use the assumption that the multiplication $\star$ on $\mathsf{N}$ is a computable function.
@@
Let us fix an enumeration of all quadrangles of the form $(m,n,D,F)$ where $m,n\in\mathbb{N}$ and $D,F\subset_{fin}\mathbb{N}$. The following procedure, denoted below by $\hat{\vartheta}(m,n,D,F)$, determines quadrangles with $F$ satisfying the condition of the proposition.
For an input $(m,n,D,F)$ let $F_{0}\subseteq F\cup D\star F$ be a set representing the $\mathsf{T}_{=}$-classes in $F\cup D\star F$. Let us fix an enumeration of the set
@@
If this is not the case we make the next step. Note that the number $\mu(\nu(F_{0}\cap F),\nu(g)\nu(F_{0}\cap F),B_{<\frac{1}{m}}(1))$ does not decrease. Indeed, since the distances are becoming smaller, new matchings can occur, but the matchings which were already found, can only increase to larger ones. We stop when the above inequality holds. Note that if in the numbered metric group $(G,d,\nu)$ the condition
@@
Fix $n_{1},n_{2}$. Let $D$ be the set $\{n_{1},n_{2}\}$. We apply the algorithm for (ii) to some $(\ell,m,n,D)$ where $\frac{2}{\ell}\leq\varepsilon$ , $m>\frac{4}{\varepsilon}$ and $\frac{3}{5}\leq\frac{n-1}{n}$. Let $F\subset\mathbb{N}$ be a set which is the output of the algorithm and let $\Sigma^{\ell}_{F}=\{(f,f^{\prime},q)\,|\,f,f^{\prime}\in F,\,q\in\mathbb{Q}^{+},d(\nu(f),\nu(f^{\prime}))\in[q,q+\frac{1}{\ell})\}$.
@@
As in Section 4 an effective Følner sequence of the group $(G,d,\nu)$ is an effective sequence $(n_{j})_{j\in\mathbb{N}}$ where each $n_{j}$ is a Gödel number of some $F_{j}$, such that $(F_{j})_{j\in\mathbb{N}}$ is a Følner sequence in $\nu(\mathbb{N})$ (i.e. $\mathbb{N}$).
@@
$(\phi(m,y)$ $\text{ is a total function})\wedge(\forall g\in\nu(\mathbb{N}))(\forall n)(\exists l)(\forall k,f)\Big(k>l\wedge(\phi(m,k)=f)$